1. TRUE/FALSE(1 mark per question, 20 marks [i] An economys income is the same as its expenditure because every transaction P) GDP is the market value of all final goods and services produced by a countrys (3] If nominal GDP is $10,000 and real GDP is $8,000, then the GDP deflator is 125. has a buyer and a seller. citizens in a given period of time. 41 Other things equal, in countries with higher levels of real GDP per person, life 15) when the consumer price index falls, the typical family has to spend fewer 6 Economists use the term inflation to describe a situation in which the economy s m When a new good is introduced, consumers have more variety from which to expectaney and literacy rates are higher than in countries with lower levels of real GDP per person. dollars to maintain the same standard of living. overall price level is rising. choose, and this in turn increases the cost of maintaining the same level of economic well-being. 18] The real interest rate is the interest rate corrected for inflation. 19] Like physical capital, human capital is a produced factor of production. oj When economists refer to investment, they mean the purchasing of stocks and [11] An increase in the demand for loanable funds increases the equilibrium interest (12) When the government budget deficit rises, national saving is reduced, interest [13) Alberta buys a paint sprayer and a lift for her car customizing shop. A bonds and other types of saving. rate and increases the equilibrium level of saving. rates rise, and investment falls. macroeconomist would refer to these purchases as investment. [14) Some degree of unemployment is inevitable in a complex economy. [1s) The natural rate of unemployment is a type of unemployment that does not g away on its own even in the long run. [16j Full-time students and homemakers are included in the Bureau of La Statistics unemployed category
Answer #1

Answer:

1. True ( every transaction has buyer and seller)

2. False ( GDP is the total market value of all final goos and services produced within the country in a given time)

3. True ( GDP deflator = (nominal GDP / real GDP)* 100 = (1000/800)*100= 125

4. False ( real GDP per person is not the true measure for life expectancy and literacy rate)

prices of geeds and cesvires) Trme om culuch to choose 1. o Trne loar [ls 13, False Iy, Tr 15, Trne Coun + ern